The radiation of energy E falls normally on a perfectly reflecting surface.What is the momentum transferred to the surface.

Grade:12th pass

2 Answers

Vikas TU
7 years ago
Momentum transferred to the surface = > mcfinal – mcinitial.
Hence,
Pi = E/c  and  Pf = E/c too.
Hence depends on the direction as,
Net Momentum = E/c – (-E/c) => 2E/c 
ankit singh
3 years ago
The radiation energy is given by. ... A radiation of energy 'E' falls normally on a perfectly reflecting surface. The momentum transferred to the surface is ... Monochromatic light of wavelength 632.8 nm is produced by a helium-neon laser.